# CAT-SAM: Conditional Tuning Network for Few-Shot Adaptation of Segmentation Anything Model

### v1

@@ 4.5 Analysis & Discussion
Prompts with Single Point. SAM provides flexibility in geometric prompts including a single foreground point. However, prompting with a single foreground point can be challenging as the single point could correspond to multiple objects [19]. Nonetheless, adaptation with few-shot target samples can alleviate this ambiguity by guiding the model to focus on the specific foreground distribution within the annotation space. We evaluate this by using the same randomly-selected single point as the geometric prompt, and Figure 3 shows the segmentation results for images from WHU, Kvasir, SBU-Shadow, and “MA. Roads” [30] on road segmentation.
@@
Figure 3: Single point to mask evaluation for one-shot adaptation.
@@
Visual Illustrations. Figure 4 shows qualitative comparisons between SAM and CAT-SAM across multiple segmentation tasks. These illustrations demonstrate how our proposed CAT-SAM remarkably enhances the mask quality of SAM under the presence of only one-shot target samples. Please refer to the appendix for more visual comparisons.
Limitations. Although CAT-SAM introduces a limited number of additional parameters, it remains computationally demanding and cannot achieve real-time speed due to its reliance on SAM, necessitating substantial GPU resources and limiting its applicability for tasks such as video processing. Additionally, its adaptation performance, particularly in highly complex domains like Sonar images with multiple classes, needs further improvements for various real-world applications especially in data-scarce scenarios.
@@ 5 Conclusion
A. Implementation Details
We provide supplementary training details beyond Section 4 of the main paper. Across all experiments, we employ the AdamW optimizer. The learning rate and weight decay are set at $1\times 10^{-3}$ and $1\times 10^{-4}$, respectively. We implement the cosine annealing strategy as the scheduler, with a minimum learning rate of $1\times 10^{-5}$. Our experiments are conducted using NVIDIA RTX A6000 GPU with 48GB of memory, where a single GPU is utilized for 1-shot adaptation and 4 GPUs for 16-shot and full-shot adaptation. In addition, for the extremely fine-grained experiments conducted on HQ-Seg44k, we adopted the same settings as in [17] for fair comparisons.
For segmentation experiments over WHU, Kvasir, SBU-Shadow, MA. Roads and HRSID, we train 1000 epochs with a batch size of 1 for 1-shot adaptation, 200 epochs with a batch size of 4 for 16-shot adaptation, and 30 epochs for the full-shot adaptation with a batch size of 4. We apply random vertical/horizontal flipping and random cropping on each dataset for data augmentation. For input geometrical prompts, we employ the same strategy as HQ-SAM [17] by randomly assigning a point, box, or coarse mask for each object.
As for segmentation over FLS and JSRT, we maintain the same training epochs and batch size except empirically train 500 epochs on 16-shot adaptation for FLS. Due to their unique geometric structures, we only adopt horizontal flipping.
@@ B. CAT-SAM across ViT Backbones
We evaluate CAT-SAM’s performance using various backbones of SAM’s image encoder. Table 9 presents the results obtained with ViT-L and ViT-H [9] image encoders for extremely fine-grained segmentation, following the same setup in Table 5 of the main paper. The results demonstrate that, in comparison to SAM [19] and HQ-SAM [17], both variants of CAT-SAM exhibit considerable adaptability when paired with different backbone image encoders. This showcases their superior efficacy in facilitating SAM adaptation. Notably, while ViT-H introduces a substantially higher number of parameters, it only yields marginal gains over ViT-L. This observation suggests that further scaling of the image encoder might not be currently advantageous, which aligns well with the findings in [19].
C. More Visual Illustrations
@@
We present additional visual illustrations showcasing the efficacy of CAT-SAM on SAM adaptation. Figure 5 shows qualitative experiments corresponding to Table 5 in the main paper, illustrating the extremely fine-grained segmentation of objects with intricate structures. These examples show SAM’s limitations clearly in generating high-quality segmentations for such objects. Though HQ-SAM achieves noticeable improvements, its segmentation remains clearly different from the ground truth. As a comparison, the two CAT-SAM variants exhibit enhanced accuracy in segmenting these challenging samples. Their superior segmentation performance underlines their effectiveness in adapting to complex structural images.
@@
Figure 6 shows qualitative experiments corresponding to Tables 6, 7, and 8 of main paper. The visual illustrations consistently corroborate the quantitative findings, demonstrating that employing more target samples consistently enhances segmentation performance across diverse and challenging RGB and non-RGB datasets. Remarkably, even with just one-shot target samples, both variants of CAT-SAM achieve remarkably high-quality target segmentation. This underscores the robustness and effectiveness of CAT-SAM in handling diverse target samples, showcasing its adaptability and generalization capabilities.
@@
D. Failure Case Analysis
Figure 7 visualizes SAM and CAT-SAM segmentation while facing challenging samples, where the experiments are conducted with 16-shot and full-shot on the FLS dataset. Though both CAT-SAM variants demonstrate great improvements over SAM under the 16-shot setup, they still exhibit clear disparities from the ground-truth segmentation. In contrast, full-shot adaptation achieves significantly better segmentation over these samples. This reinforces the points we highlighted in the limitations part of this study: more research is needed for accurate and reliable segmentation especially while handling data-scarcity scenarios in various real-world tasks.

### v3

@@ 4.5 Analysis & Discussion
Figure 3: Single point to mask evaluation for one-shot adaptation.
Prompts with Single Point. SAM provides flexibility in geometric prompts including a single foreground point. However, prompting with a single foreground point can be challenging as the single point could correspond to multiple objects [23]. Nonetheless, adaptation with few-shot target samples can alleviate this ambiguity by guiding the model to focus on the specific foreground distribution within the annotation space. We evaluate this by using the same randomly-selected single point as the geometric prompt, and Fig. 3 shows the segmentation results for images from WHU, Kvasir, SBU-Shadow, and MA. Roads [32] on road segmentation.
@@
Visual Illustrations. Fig. 4 shows qualitative comparisons between SAM and CAT-SAM across multiple segmentation tasks. These illustrations demonstrate how our proposed CAT-SAM remarkably enhances the mask quality of SAM under the presence of only one-shot target samples. Please refer to the appendix for more visual comparisons.
@@ 5 Conclusion
Appendix 0.A Implementation Details
We provide supplementary training details beyond Section 4 of the submitted manuscript. Across all experiments, we employ the AdamW optimizer. The learning rate and weight decay are set at $1\times 10^{-3}$ and $1\times 10^{-4}$, respectively. We implement the cosine annealing strategy as the scheduler, with a minimum learning rate of $1\times 10^{-5}$. Our experiments are conducted using NVIDIA RTX A6000 GPU with 48GB of memory, where a single GPU is utilized for 1-shot adaptation and 4 GPUs for 16-shot and full-shot adaptation. In addition, for the extremely fine-grained experiments conducted on HQ-Seg44k, we adopted the same settings as in [21] for fair comparisons.
For segmentation experiments over WHU, Kvasir, SBU-Shadow, MA. Roads and HRSID, we train 1000 epochs with a batch size of 1 for 1-shot adaptation, 200 epochs with a batch size of 4 for 16-shot adaptation, and 30 epochs for the full-shot adaptation with a batch size of 4. Table 4 at the end of this appendix lists the names of the used few-shot training samples. We apply random vertical/horizontal flipping and random cropping on each dataset for data augmentation. For input geometrical prompts, we employ the same strategy as HQ-SAM [21] by randomly assigning a point, box, or coarse mask for each object.
We adopted the same training epochs and batch sizes for segmentation experiments on FLS and JSRT, except that we trained 500 epochs on 16-shot adaptation for FLS. For data augmentation, we adopted horizontal flipping only due to the unique geometric structures of the two datasets.
Appendix 0.B Mask Decoder Details
As depicted in Fig. 5, we freeze the entire decoder of SAM and introduce a learnable CAT-token ($1\times 256$). This token is concatenated with the output tokens and prompt tokens of the original SAM, forming the input to the mask decoder of CAT-SAM. Following the same pipeline of SAM, the input passes through two decoder layers, and the updated CAT-Token is then employed to generate dynamic weights for a newly introduced three-layer MLP. Additionally, following [21], we upsample and fuse the output image features from the 6th and the last layers of the image encoder. Ultimately, this fused feature is combined with SAM’s mask decoder features, and further with the output of the three-layer MLP (via dot product) to produce the target mask.
Figure 5: Details of the mask decoder in CAT-SAM. Both CAT-SAM variants share the same decoder tuning structures.
@@ Appendix 0.C CAT-SAM across ViT Backbones
We evaluate CAT-SAM’s performance using various backbones of SAM’s image encoder. Tab. 8 presents experiments with ViT-L and ViT-H [9] image encoders on extremely fine-grained segmentation, following the same setup in Table 6 of our submitted manuscript. The results show that, in comparison to SAM [23] and HQ-SAM [21], both CAT-SAM variants exhibit considerable adaptability when paired with different backbone image encoders. This showcases their superior efficacy in facilitating SAM adaptation. Notably, while ViT-H introduces a substantially higher number of parameters, it only yields marginal performance gains over ViT-L. This observation suggests that scaling up the image encoder further might not give much edge, which aligns well with the findings in [23].
Appendix 0.D More Experimental Results
Per-Class Performance on FLS. Table 9 details the per-class performance from Table 7 of the main paper for marine debris segmentation using Sonar images on the FLS dataset [40]. We compare SAM and CAT-SAM variants across 11 classes and their average segmentation results. CAT-SAM variants consistently exhibit superior adaptation effects, with performance improving as the number of samples increases.
Table 9: Few-shot adaptation with varying number of training samples for non-RGB datasets, including FLS [40] with Sonar images for marine debris segmentation. 1-shot adaptation unattainable due to the absence of images containing objects from all 11 classes.
#Samples Methods FLS [40]
Bott. Can Chain Drin. Hook Prop. Sham. Stan. Tire Valve Wall mIoU
None SAM [23] 76.0 81.4 38.4 78.5 47.9 71.6 81.7 85.7 69.3 54.0 82.7 69.7
1-shot CAT-SAM-T
CAT-SAM-A
16-shot CAT-SAM-T 78.1 76.7 46.4 82.3 49.5 73.5 82.8 85.6 83.2 61.8 85.0 73.2
CAT-SAM-A 72.4 72.6 47.9 78.9 52.7 69.6 81.6 87.4 73.8 62.0 86.3 71.4
Full-shot CAT-SAM-T 86.3 87.9 66.7 83.2 67.0 79.5 87.0 88.7 90.9 70.3 91.3 81.7
CAT-SAM-A 85.5 88.1 67.1 83.5 67.6 80.2 86.5 89.8 91.2 69.9 92.1 82.0
Table 10: Comparison of adaptive segmentation performance on challenging non-rgb datasets: JSRT [39] for chest X-ray segmentation, FLS [40] with Sonar images for marine debris segmentation, and HRSID [45] for ship instance segmentation with SAR images. The baseline is SAM without adaptation. We report 1-shot/16-shots/full-shots adaptation results.
     Model      JSRT (mIoU)      FLS (mIoU)      HRSID (AP)
     LoRA[18]      88.1/94.0/94.2      -/70.1/80.2      41.0/44.7/50.8
     HQ-SAM[24]      89.0/92.1/93.1      -/70.2/79.5      40.1/41.4/48.1
     CAT-SAM-T      93.0/94.2/94.4      -/73.2/81.7      46.0/46.2/51.4
     CAT-SAM-A      92.6/93.5/94.6      -/71.4/82.0      44.9/45.7/52.9
Comparison with the SoTA for Non-RGB Domains. In Table 10, we benchmark CAT-SAM against the leading methods from each group in Table 4 of the main paper, specifically LoRA [16] and HQ-SAM [21], across non-RGB domains including JSRT, FLS, and HRSID. Experiments conducted with 1-shot, 16-shot, and full-shot adaptations consistently demonstrate that CAT-SAM outperforms these methods across all three datasets, particularly with fewer training samples, supporting our conclusions in the main paper.
Scaled Parameters. We investigate the impact of increasing the number of tokens in the image encoder of CAT-SAM-T (denoted as $P_{i}$) on adaptation effects. Table 11 presents the results of 1-shot adaptation with varying token and parameter counts. Doubling the tokens from 1 to 2 results in marginal performance gains across the three datasets. However, further increases in the token count lead to a gradual decline in performance, likely due to the increased parameters requiring more training data for effective tuning. This underscores the effectiveness of using fewer tokens and parameters for few-shot adaptation tasks.
Table 11: Segmentation of CAT-SAM-T with varing number of tokens in the image encoder for 1-shot adaptation.
Token number 1 2 4 8 16 32
#Paras 3.3M 5.3M 9.3M 17.3M 33.2M 65.1M
WHU 86.8 87.4 85.7 84.3 82.4 80.2
Kvasir 83.4 85.2 81.4 81.3 79.1 73.9
SBU 78.0 79.0 76.6 75.3 70.3 67.9
Appendix 0.E More Visual Illustrations
@@
We present additional visual illustrations showcasing the efficacy of CAT-SAM on SAM adaptation. Fig. 6 shows qualitative experiments corresponding to Table 6 in our submitted manuscript, illustrating the extremely fine-grained segmentation of objects with intricate structures. These examples show SAM’s limitations clearly in generating high-quality segmentations for such objects. Though HQ-SAM achieves noticeable improvements, its segmentation remains clearly different from the ground truth. As a comparison, the two CAT-SAM variants exhibit much better accuracy in segmenting these challenging samples. The superior segmentation performance of the two CAT-SAM variants underlines their effectiveness in adapting to complex structural images.
@@
Fig. 7 shows qualitative experiments corresponding to Tables 5, 7, and 8 of our submitted manuscript. The visual illustrations consistently corroborate the quantitative findings, demonstrating that employing more target samples consistently enhances segmentation performance across diverse and challenging RGB and non-RGB datasets. Remarkably, even with just one-shot target samples, both variants of CAT-SAM achieve remarkably high-quality target segmentation. This underscores the robustness and effectiveness of CAT-SAM in handling diverse target samples, showcasing its adaptability and generalization capabilities.
Appendix 0.F Robustness Analysis
We examine the robustness of CAT-SAM by evaluating it with diverse target samples under the challenging 1-shot adaptation setup. Specifically, we randomly select three distinct training images from each of the WHU, Kvasir, and SBU-Shadow datasets and train three 1-shot adaptation models, each following the same setup as detailed in Table 4 of the manuscript. Subsequently, we report the averaged segmentation performance over the testing set for the three models, along with upper and lower deviations. Tab. 12 summarizes the experimental results. It is evident that across all three datasets, both variants of CAT-SAM consistently demonstrate notable 1-shot adaptation effects with varying training samples, demonstrating the superior robustness of our CAT-SAM.
Table 12: 1-shot adaption of CAT-SAM variants across challenging true color image datasets. Both CAT-SAM variants achieve superior and robust adaptation results when trained with different training samples.
Methods WHU Kvasir SBU-Shadow
SAM [23] (baseline) 43.5 79.0 62.4
CAT-SAM-T (Ours) 87.3±0.5 83.2±0.5 79.0±1.0
CAT-SAM-A (Ours) 88.5±0.4 85.7±0.5 84.1±2.2
Appendix 0.G Failure Case Analysis
Fig. 8 visualizes SAM and CAT-SAM segmentation while facing challenging samples, where the experiments are conducted with 16-shot and full-shot on the FLS dataset. Though both CAT-SAM variants demonstrate great improvements over SAM under the 16-shot setup, they still exhibit clear disparities from the ground-truth segmentation. In contrast, full-shot adaptation achieves significantly better segmentation over these samples. This is consistent with the points we stated in the limitations part of the manuscript: more research is needed for accurate and reliable segmentation especially while handling data-scarcity scenarios in challenging real-world tasks.
@@ Appendix 0.G Failure Case Analysis
Appendix 0.H Limitations
Although CAT-SAM introduces a limited number of additional parameters, it remains computationally demanding and cannot achieve real-time speed due to its reliance on SAM, necessitating substantial GPU resources and limiting its applicability for tasks such as video processing. Additionally, its adaptation performance, particularly in highly complex domains like Sonar images with multiple classes, needs further improvements for various real-world applications especially in data-scarce scenarios.
Appendix 0.I More Experimental Details
LABEL:tab:filenames lists the file names of training samples employed in both 1-shot and 16-shot adaptation experiments as described in the manuscript.
Table 13: File names for 1-shot and 16-shot training.
Dataset File Names
1-shot
Kvasir cju0roawvklrq0799vmjorwfv
WHU 2432
SBU lssd1409
JSRT case164
M Road 23879065_15
HRSID P0003_600_1400_9189_9989
16-shot
Kvasir cju7dn24o296i09871qfxb8s2,
cju83mki1jv5w0817kubxm31r,
cju6vucxvvlda0755j7msqnya,
cju43kj2pm34f0850l28ahpni,
cju2x7vw87mu30878hye2ca0m,
cju5ddda9bkkt0850enzwatb1,
cju77vvcwzcm50850lzoykuva,
cju3128yi0rpu0988o4oo5n8n,
cju8b7aqtr4a00987coba14b7,
cjz14qsk2wci60794un9ozwmw,
cju0s690hkp960855tjuaqvv0,
cju888fr7nveu0818r9uwtiit,
cju8b542nr81x0871uxnkm9ih,
cju2qtee81yd708787bsjr75d,
cju83h9ysjwe808716nt35oah,
cju0roawvklrq0799vmjorwfv
WHU 168, 515, 807, 1176, 1552, 1901, 2432, 2679,
3130, 3486, 3678, 3939, 4263, 4335, 4640, 4725
SBU lssd3356, lssd1409, lssd1090, lssd3732,
lssd2011, lssd1901, lssd1821, lssd1512,
lssd3713, lssd1376, lssd3493, lssd3729,
lssd610, lssd3008, lssd1319, lssd1109
JSRT case007, case009, case023, case027,
case029, case036, case058, case063,
case071, case109, case140, case152,
case164, case174, case189, case190
M Road 10528660_15, 10528735_15, 15928885_15, 20128945_15,
20578960_15, 21928990_15, 22378945_15, 22979005_15,
23278960_15, 23428960_15, 23728540_15, 23879065_15,
24178525_15, 24628855_15, 24629290_15, 26128780_15
FLS marine-debris-aris3k-53, marine-debris-aris3k-295,
marine-debris-aris3k-298, marine-debris-aris3k-322,
marine-debris-aris3k-365, marine-debris-aris3k-391,
marine-debris-aris3k-744, marine-debris-aris3k-1039,
marine-debris-aris3k-1307, marine-debris-aris3k-1324,
marine-debris-aris3k-1359, marine-debris-aris3k-1393,
marine-debris-aris3k-1457, marine-debris-aris3k-1774,
marine-debris-aris3k-1816, marine-debris-aris3k-1817
HRSID P0081_2400_3200_4200_5000,
P0012_5400_6200_7200_8000,
P0105_3600_4400_16200_17000,
P0010_2400_3200_4800_5600,
P0076_0_800_4200_5000,
P0064_1800_2600_5400_6200,
P0060_1200_2000_6000_6800,
P0056_1200_2000_5400_6200,
P0067_600_1400_6000_6800,
P0002_1200_2000_8400_9200,
P0027_1800_2600_8289_9089,
P0109_3000_3800_7200_8000,
P0108_2400_3200_7200_8000,
P0109_3000_3800_9600_10400,
P0105_3600_4400_15600_16400,
P0028_600_1400_7200_8000